OpenStudy (anonymous):

People who give blood give about a pint at a time. How many cc's (cubic centimeters) is this?

OpenStudy (anonymous):

1 liquid gal = 231 cubic inches 1 liquid qt= 1/4 liquid gal = 32 fl oz. 1 liquid pt. = 16 fl oz 1 liter=1000 cubic centimeters 1 mL=1/1000 liters